OpenStudy (anonymous):

what's the answer to 6^x=24 ?

Parth (parthkohli):

Take the \(\log\) on both sides. Okay?

Parth (parthkohli):

Also, do remember that,\[a^x = b\implies \log_a b = x\]

Parth (parthkohli):

So,\[\log_{6}24 = x\]\[{\log{24} \over \log 6} = x \]Input that into a calculator!

OpenStudy (anonymous):